OpenStudy (anonymous):

Please help! Simplify the radical expression: (9-sqrt45)/(3)

OpenStudy (anonymous):

(9-sqrt45)/(3) =(9-3*sqrt5)/(3) =(3-sqrt5)

OpenStudy (anonymous):

Thanks! @bhaskarbabu